Blast it! said David Wilson, president of Teledex Company. Weve just lost the bid on the Koopers job by $3,000. It seemsBlast it! said David Wilson, president of Teledex Company. Weve just lost the bid on the Koopers job by $3,000. It seems
Answer #1

Solution 1a:

Overhead rate = 894600/639000 = 140%

Solution 1b:

Total manufacturing Overhead applied = 13500*140% = 18900

Solution 2a and 2b:

Computation of overhead rate for Each department
Particulars Fabricating Machining Assembly
Overhead cost of deparment $3,72,750 $4,26,000 $95,850
/Total direct labor Cost $2,13,000 $1,06,500 $3,19,500
Overhead rate of department 175% 400% 30%
Computation of Manufacturing Overhead applied to Koopers Job
Particulars Fabricating Machining Assembly
Direct Labor $5,400 $600 $7,500
Overhead rate of department 175% 400% 30%
Manufacturing overhead applied $9,450 $2,400 $2,250
Total Overhead applied $14,100

Solution 4a and 4b:

Computation of Bid price on Koopers Job (Using Plantwide Overhead rate)
Particulars Amount
Total Direct material $7,300
Total Direct Labor $13,500
Total manufacturing Overhead applied $18,900
Total manufacturing cost $39,700
Bid Price ($39700*150%) $59,550
Computation of Bid price on Koopers Job (Using Departmental Overhead rate)
Particulars Amount
Total Direct material $7,300
Total Direct Labor $13,500
Total manufacturing Overhead applied $14,100
Total manufacturing cost $34,900
Bid Price ($34900*150%) $52,350